### v4

# Here the subspaces are defined by

1. Introduction
A classical theorem of Hilbert [11] asserts that if is a complete surface of constant Gaussian curvature in then The analogous statement for hypersurfaces with constant scalar curvature in was shown to be false by Okayasu [14]. The main purpose of the present paper is to study this question for complete submanifolds of space forms under certain additional conditions.
Let denote the complete simply-connected Riemannian manifold of dimension and constant sectional curvature Let be an isometric immersion with flat normal bundle whose second fundamental form we denote by Then for any there exists a positive integer and unique pairwise distinct normal vectors at called the principal normal vectors of at such that
Here the subspaces are defined by
We say that is a proper immersion if the integer is independent of In this case, it is known [16] that each is a smooth normal field on and that the distributions are smooth. For convenience, we say that has type for some integer if for each In this paper we consider certain isometric immersions into space forms which have type
Examples. Suppose that is an isometric immersion with flat normal bundle that is proper and has no umbilic points.
(a) Assume that is conformally flat and By a theorem of Cartan [1], has a principal curvature of multiplicity at each point. Hence has type
(b) Assume that is conformally flat as before and Although is not necessarily of type by combining a result of Moore [15] on the existence of a principal normal vector field of high multiplicity with that of Dajczer, Onti and Vlachos ([6], Th. ), we infer that the possible types of are limited when the codimension is small. In particular, when and the only possible type of other than the type is
Our first result is the following theorem.
Theorem 1.1.
Let be a complete Riemannian manifold with constant scalar curvature and be an isometric immersion with flat normal bundle which is proper and has type Then the following holds.
(i) if The inequality is strict if is compact.
(ii) if
(iii) if The inequality is strict if is compact.
The inequalities in Theorem 1.1 are optimal. Moreover, Theorem 1.1 is global in the sense that the completeness assumption on cannot be dropped. See for details.
Remark. Parts (ii) and (iii) of Theorem 1.1 extend some results of Cheng ([2], Th. ), Hu and Zhai ([12], Th. ) (also see Leite [13]) on hypersurfaces to submanifolds with flat normal bundle.
The main idea in the proof of Theorem 1.1 is to work in principal coordinate systems and use a convexity argument. The same method also yields the following result which is of a local character.
Theorem 1.2.
Let be a Riemannian manifold with constant scalar curvature and be an isometric immersion with flat normal bundle which is proper and has type Assume that the mean curvature field of is parallel. Then the sectional curvature of satisfies
Remark. We note that the flat normal bundle assumption in Theorem 1.2 is redundant when the codimension and is not minimal ([9], Th. ).
Combining Theorem 1.2 with a result of Erbacher ([9], Th. ) we readily obtain the following classification.
Corollary 1.3.
Let be a complete Riemannian manifold with constant scalar curvature and be an isometric immersion with flat normal bundle which is proper and has type Assume that the mean curvature field of is parallel. Then must be of the form or if and must be of the form if The corresponding local result holds if the completeness assumption on is dropped.
2. Preliminaries
Assume that and are as in Theorem 1.1. In this case, the distributions are integrable. Arguing as in [6] we see that each point admits a principal coordinate system such that and on In such a coordinate system the first and the second fundamental forms of the immersion are given respectively by
and
for certain positive functions defined on
The Gauss equation for an isometric immersion in applied to the orthonormal set
yields
and
Here denotes the sectional curvature of along the plane spanned by the vectors and
The Codazzi equations ([7], p. 36)
and
now take the forms
for any
Let Since has scalar curvature we obtain
For later use in the proof of Theorem 1.1, we note that in the special case when for each we may assume that Making the change of variable
we see that the first fundamental form of in the resulting principal coordinate system is given by
Finally, we assume that is and connected.
3. A Scalar Curvature Inequality
This section is devoted to the proof of Theorem 1.1.
Define by
Since is constant by our assumption, in any principal coordinate system in a straightforward computation using (2.4) and (2.6) gives
Moreover, from (2.3) it follows that
On the other hand, using (2.3) and (2.5) we obtain
If for some then Therefore (2.6) gives
From (3.1) and it follows that for every Thus has constant sectional curvature and, since it follows that in this case.
Therefore we may assume that for any From (2.5) we conclude for each As noted in it follows that each point of admits a principal coordinate system in which the first fundamental form of the immersion is given by
In the remainder of the proof we will only work with principal coordinate systems on that are of the above form.
Since we have Let and A straightforward computation ([7], p. 20) gives
and
for any
Let denote the curvature tensor on We compute
Thus
Fix a point and a principal coordinate system around as above. Let denote the geodesic with initial data and Since the integral curves of are geodesics parametrized by arc-length it follows from uniqueness that
whenever If denotes another principal coordinate system that also contains for some then, since and is -dimensional, we must have Replacing by if necessary, we can assume that In this way we obtain a covering of by principal coordinate charts which satisfy the condition
whenever
Let Using (2.6) we obtain
Assume that the function is positive for the moment. In this case, we define a function by
Then is positive and satisfies for each Working in a principal coordinate system we obtain
Here β² denotes derivative with respect to the arc-length parameter Using
and (3.2) we obtain
Thus
Now we consider the three cases separately.
(i) Let We assume that and obtain a contradiction. Using (2.6) we obtain
Also, since from (3.4) it follows that Since is strictly convex on we have This contradicts for all Hence
Now assume that is compact and Using (2.6) we obtain Assume that attains minimum at We claim that If then from (3.3) we obtain The equation
which is obtained from (3.3) by differentiation and the condition together imply that Using and (3.2) we conclude
Hence and which contradicts the assumption. Therefore From (3.3) and we conclude that vanishes on From (2.6) we get and it follows that must be flat. Since vanishes identically, has constant index of relative nullity From Hartmanβs theorem [10] it follows that must be an -cylinder and contains a line. This contradicts the compactness of
(ii) Let Suppose From (2.6) we have
Moreover, (2.6) implies that or From (3.4) it follows that is convex, and strictly convex if As in part (i) we obtain a contradiction if or if and one of the limits is infinite. Otherwise must be a constant function and and hence must be constant along Since it follows that must be a constant function on From (3.3) we obtain for each Hence we obtain using (3.2). Since using (2.6) we obtain This gives which is a contradiction.
(iii) Let Suppose Since from (2.6) we obtain Using (2.6) once more we get or since by assumption. Rest of the proof is analogous to part (i).
In the case of equality, we have so that for every However, if is compact then the argument in part (i) shows that there exists a point such that Hence the inequality must be strict if is compact.
β
Remarks. Below we give examples to show that Theorem 1.1 is global and the bounds obtained are optimal. These examples are constructed using the theory of rotation hypersurfaces in space forms [5]. Below we will follow the notation and terminology of this paper. Recall that For simplicity, we consider the case
To see that the completeness assumption on cannot be dropped, let be any orthogonal parametrization of the unit sphere in For sufficiently small define
Then Consider the rotation hypersurface in given by
and satisfying the arc-length parametrization condition By a straightforward computation using formulas for the principal curvatures of a rotation hypersurface ([5], p. 693) and (2.6) we obtain
It may be checked that the arc-length parametrization condition cannot be continued for all Similar examples can be constructed in any dimension and any space form by solving a certain ODE (See Leite [13]).
The bounds in Theorem 1.1 are also optimal. For the circular cylinder satisfies Suppose that and Let be as in the previous example. For any real number let Then the rotation hypersurface in which is given by
and satisfying the arc-length parametrization condition on the meridian curve
satisfies We note that the Clifford minimal hypersurface in corresponds to the choice Finally, for and let Then the parabolic rotation hypersurface in obtained from the profile function satisfies The last two statements are easily verified using a simple computation as in the first example.
4. A Sectional Curvature Inequality
In this section we give proof of Theorem 1.2.
As with the proof of Theorem 1.1 we work in a principal coordinate system on We claim that vanishes identically for any Otherwise, by shrinking if necessary, we may assume that does not vanish on Let be the mean curvature field of Since and from (2.3) and (2.5) it follows that
and
Thus for each From (2.4) we conclude that
Since has flat normal bundle using (2.3) we obtain
Substituting for from (2.4) and simplifying we obtain
From (4.1) it follows that
Combining this with equation (4.3) we obtain
for any Hence we may assume that
for some positive functions and Differentiating (2.6) with respect to we get
Substituting from (4.2) in the above equation gives
Since we have so that vanishes identically and hence From this and we conclude as in the proof of Theorem 1.1. Since the curvature operator of is diagonalized by the orthonormal basis we conclude that the sectional curvatures of lie between and