### v5

# Some curvature inequalities for submanifolds

Abstract.
In this paper, we first study isometric immersions into space forms with flat normal bundle and constant scalar curvature Under a suitable multiplicity condition on the second fundamental form of we prove the global result that if and if We then obtain a classification for submanifolds with flat normal bundle having constant scalar and mean curvatures. Finally, in the hypersurface case, analogous results are proved for conformally flat hypersurfaces with constant higher order mean curvatures.
Key words and phrases:
Isometric immersion, Flat normal bundle, Second fundamental form, Higher order mean curvature, Conformally flat manifold2010 Mathematics Subject Classification:
Primary 53C40; Secondary 53C421. Introduction
A classical theorem of Hilbert [13] asserts that if is a complete surface in with constant Gaussian curvature then Cartan [2, 3] showed that a Riemannian manifold with dimension and having constant negative sectional curvature cannot be isometrically immersed in for purely local reasons. On the other hand, he constructed many local examples of submanifolds of having constant negative sectional curvature and proved that any such example must necessarily have flat normal bundle. Since then it had been a major open problem in submanifold geometry whether a complete Riemannian manifold with constant negative sectional curvature admits an isometric immersion into In a recent preprint Moore [18] has answered this question in the negative, thereby extending Hilbert’s theorem to higher dimension.
It is natural to ask whether similar results hold for the scalar curvature. However, Okayasu [20] constructed the first examples of complete hypersurfaces in with constant negative scalar curvature. A primary purpose of this paper is to present a variant of Hilbert’s theorem for the scalar curvature in general codimension and space form under an additional condition which we describe below.
Let denote the complete simply-connected Riemannian manifold of dimension and constant sectional curvature Let be an isometric immersion with flat normal bundle whose second fundamental form we denote by It is well-known ([9], p. 35) that for each there exists a positive integer and unique pairwise distinct normal vectors at called the principal normal vectors of at such that
where the subspaces are defined by
We call a proper immersion if the integer is independent of In this case, it is known [22] that each defines a smooth normal field on and the distributions are smooth. For convenience, we say that has type for some integer if for each
In this paper we will be primarily concerned with isometric immersions into space forms that have type In the hypersurface case, this condition implies that is conformally flat and, conversely, if is conformally flat then has type provided it has no umbilical points. This follows from a well-known result of Cartan [4]. In general codimension, Lobos and Tojeiro ([15], proposition 2.5) showed that if and has type with linearly independent principal normal vector fields, then where is a quasiumbilical hypersurface (i.e. it has a principal curvature of multiplicity at least at every point) and is an isometric immersion with no totally-geodesic points on an open set In addition, on the open set of non-umbilical points of each leaf of the umbilical foliation of is mapped into a relative nullity leaf of Conversely, any such isometric immersion with free of umbilical points, has type with linearly independent principal normal vector fields.
Remark 1.1.
Suppose is conformally flat and and is an isometric immersion that is proper and has type different from Although is not necessarily of type in this case, combining a result of Moore [17] on the existence of a principal normal vector such that for each with a result of Dajczer, Onti and Vlachos ([7], theorem ), we infer that the possible types of are limited when the codimension is small. In particular, the following important special case in codimension two should be noted:
Let be conformally flat and be a proper isometric immersion that is not totally-umbilical. If and then the only possible type of other than is
If is umbilical at a point (i.e. has precisely one principal normal vector at ), then it follows trivially from the Gauss equation that the scalar curvature of at is greater than or equal to Our first result is of a global nature.
Theorem 1.2.
Let be a complete Riemannian manifold with constant scalar curvature and be an isometric immersion with flat normal bundle which is proper and has type Then the following holds.
(i) if The inequality is strict if is compact.
(ii) if
(iii) if The inequality is strict if is compact.
The inequalities in Theorem 1.2 are optimal. We emphasize that Theorem 1.2 is global and the completeness assumption on cannot be dropped. See for examples.
Remark 1.3.
In the hypersurface case, Theorem 1.2 is a consequence of the theory of rotation hypersurfaces in space forms due to do-Carmo and Dajczer [6], and the study of rotation hypersurfaces of constant scalar curvature due to Leite [14]. However, our proof of Theorem 1.2 is direct and is independent of the theory of rotation hypersurfaces or [14]. Moreover, our method can be readily adopted to the setting of higher order mean curvatures of a hypersurface. See
The main idea behind the proof of Theorem 1.2 is to work with principal coordinate systems (see ) and use a convexity argument. The same method also yields the following result which is of a local character.
Theorem 1.4.
Let be a Riemannian manifold with constant scalar curvature and be an isometric immersion with flat normal bundle which is proper and has type Assume that the mean curvature field of is parallel. Then the sectional curvature of satisfies
Remark 1.5.
Combining Theorem 1.4 with a result of Erbacher ([10], theorem ) we immediately obtain the following classification for type isometric immersions with constant scalar and mean curvatures (cf. [8], theorem 1; also [25], [5], [23]).
Corollary 1.6.
Let be a complete Riemannian manifold with constant scalar curvature and be an isometric immersion with flat normal bundle which is proper and has type Assume that the mean curvature field of is parallel. Then must be of the form or if and must be of the form if The corresponding local result holds if the completeness assumption on is dropped.
In we specialize to hypersurfaces and derive an analogue of Theorem 1.2 for the higher order mean curvatures of a hypersurface. Let be an orientable hypersurface and be any integer. The -th higher order mean curvature of is defined by
where denotes the set of principal curvatures of with respect to a chosen smooth unit normal field on In particular, is the mean curvature of and is the Gauss-Kronecker curvature of It is well-known that is intrinsic when is even, and that is intrinsic.
The method of proof of Theorem 1.2 can be readily adopted to prove the following result.
Theorem 1.7.
Assume that is a complete conformally flat orientable hypersurface with constant -th higher order mean curvature If is even and then
In Theorem 1.7, suppose is compact. If there exists a point such that the principal curvatures of at are all strictly positive with respect to a suitably chosen normal. Consequently, for each even Same conclusion holds for but not for
Finally, we show that our earlier result ([11], theorem 1.1) extends to a conformally flat hypersurface with constant -th higher order mean curvature for any
Proposition 1.8.
Let be a connected conformally flat orientable hypersurface of a space form Assume that has constant -th higher order mean curvature for some integer Then the following holds.
(i) Suppose Then either is totally-umbilical or it has no umbilical points and is an open subset of a rotation hypersurface in
(ii) Suppose Then either has constant sectional curvature or it has no umbilical points and is an open subset of a rotation hypersurface in
2. Preliminaries
Assume that and are as in Theorem 1.2. Let and denote the principal normal vectors at in with multiplicities and respectively. In this case, it is well-known that the distributions and are integrable and the immersion is locally holonomic. Thus each point in admits a principal coordinate system such that and on In these coordinates, the first and the second fundamental forms of are given respectively by
and
where and are some positive functions defined on
Applying the Gauss equation to the orthonormal set
and making use of the expression for the second fundamental form of given above yields
and
Here denotes the sectional curvature of along the -plane spanned by the vectors and
For any vector bundle let denote the set of all smooth local sections of In the current setting, the Codazzi equations take the form ([9], p. 36, equation (1.41))
and
Substituting coordinate vector fields for and in these equations and using the expression for the first fundamental form of given above we obtain
and
where
3. A Scalar Curvature Inequality
This section is devoted to the proof of Theorem 1.2. We begin with a few preparatory lemmas.
Define two functions by
and
Let be any connected principal coordinate chart in Since is constant by assumption, using (2.4) and (2.6) we compute
Therefore
Similarly, from (2.3) it follows that
Thus
Recall from that and
Lemma 3.1.
If for some then has constant sectional curvature and
Proof. Since we have Therefore (2.6) gives From (3.1) and (3.2) we get for every and consequently for every by covering by a collection of overlapping principal coordinate charts. Since has flat normal bundle and the curvature operator of is diagonalized by the orthonormal basis it follows that has constant sectional curvature ([21], p. 63). Further, since we conclude in this case. ∎
In view of Lemma 3.1 we may assume that for every From (3.3) for each After fixing a point and making the change of variable
we may assume that the first fundamental form of in the resulting principal coordinates is given by
In the remainder of this proof we assume that this reduction has already been made and only work with principal coordinate systems on that are of this form.
Lemma 3.2.
We have for any with
Proof. This is standard. See, for example, ([6], p. 20). ∎
Fix a point and a principal coordinate system around Let denote the geodesic with initial data and Note that the geodesic is defined over since is geodesically complete. Since the integral curves of are geodesics parametrized by arc-length, it follows from uniqueness that
whenever If denotes another principal coordinate system that also contains for some then, since and is -dimensional, we must have Replacing by if necessary, we can assume that In this way we obtain a covering of by principal coordinate charts which satisfy the condition
whenever
Assuming that the function is positive for the moment, we define a function by
Lemma 3.3.
The function satisfies the equation for each
Proof. Working in a principal coordinate system we obtain
Here ′ denotes derivative with respect to the arc-length parameter of Using
and Lemma 3.2 we obtain
and
Therefore
Now we consider the three cases separately and complete the proof of Theorem 1.2.
(i) Let We assume that and obtain a contradiction. Using (2.6) we obtain
Also, since from Lemma 3.3 it follows that Since is strictly convex on we have This contradicts for all Hence
Now assume that is compact and Using (2.6) we obtain Assume that attains minimum at We claim that If then from (3.1) we obtain The equation
which is obtained from (3.1) by differentiation, and the condition together imply that Using (2.1) and Lemma 3.2 we conclude
Hence and which contradicts the assumption. Therefore From (3.1) and we conclude that vanishes on From (2.6) we get and it follows that must be flat. Since vanishes identically, has constant index of relative nullity From Hartman’s theorem [12] must be an -cylinder and contains a line. This contradicts the compactness of
(ii) Let Suppose From (2.6) we have
Moreover, (2.6) implies that or From Lemma 3.2 it follows that is convex, and strictly convex if As in part (i) we obtain a contradiction if or if and one of the limits is infinite. Otherwise must be a constant function and and hence must be constant along Since it follows that must be a constant function on From (3.1) we obtain for each Hence using Lemma 3.2. Since using (2.6) we obtain This gives which is a contradiction.
(iii) Let Suppose Since from (2.6) we obtain
Using (2.6) once more we get or since by assumption. Rest of the proof is analogous to part (i).
In the case of equality, we have so that for every However, if is compact then the argument in part (i) shows that there exists a point such that Hence the inequality must be strict if is compact. ∎
Example 3.4.
Below we give examples to show that Theorem 1.2 is global and the bounds obtained are optimal. These explicit examples are constructed using the theory of rotation hypersurfaces in space forms [6]. Below we will follow the notation and terminology of this paper. Recall that For simplicity, we consider the case
To see that the completeness assumption on cannot be dropped, let
be any orthogonal parametrization of the unit sphere in For sufficiently small we define
Then Consider the rotation hypersurface in given by
and satisfying the arc-length parametrization condition By a straightforward computation using formulas for the principal curvatures of a rotation hypersurface ([6], p. 693) and (2.6) we obtain
It may be checked that the arc-length parametrization condition cannot be continued for all Similar examples can be constructed in any dimension and any space form by solving a certain ODE (See Leite [14]).
The bounds in Theorem 1.2 are also optimal. For the circular cylinder satisfies Suppose that and Let be as in the previous example. For any real number let Then the rotation hypersurface in which is given by
and satisfying the arc-length parametrization condition on the meridian curve
satisfies We note that the Clifford minimal hypersurface in corresponds to the choice Finally, for and let Then the parabolic rotation hypersurface in obtained from the profile function satisfies The last two statements are easily verified using a simple computation as in the first example.
4. A Sectional Curvature Inequality
In this section we prove Theorem 1.4.
As in the proof of Theorem 1.2 we work in a principal coordinate system on We claim that vanishes identically for every Otherwise, by shrinking if necessary, we may assume that does not vanish on Let be the mean curvature field of Since and from (2.3) and (2.5) it follows that
and
Thus for each From (2.4) we conclude that
Since has flat normal bundle, using (2.3) we obtain
From (4.1) it follows that
for some positive functions and Differentiating (2.6) with respect to we obtain
Substituting from (4.2) in the above equation gives
5. Hypersurfaces and higher order mean curvature
Let be a hypersurface and be any integer. The -th higher order mean curvature of is defined by
where denotes the set of principal curvatures of with respect to a smooth (local) unit normal field. In particular, is the mean curvature of and is the Gauss-Kronecker curvature of the hypersurface. It is well-known that is intrinsic when is even, and that is intrinsic. Hypersurfaces with constant higher order mean curvatures are extensively studied (cf. [1], [19], [16], [24]).
Proof of Theorem 1.7. We only need to modify the proof of Theorem 1.2 slightly. If is umbilical at some point then the conclusion is trivial. Otherwise, since is conformally flat and by the aforementioned result of Cartan [4] we may assume that the principal curvatures of satisfy the relation and where and denote the distinct principal curvatures of with respect to a smooth unit normal field on Let and Then and are the principal normal vector fields of and has type From here onwards we proceed as in the proof of Theorem 1.2. The Codazzi equations can be written in terms of the principal curvatures as
for every
By definition
Let
Define functions by
and
Note that is well-defined since is even.
Let be any principal coordinate system in Since is constant by assumption, from (5.1) and (5.4) it follows that
Thus
If there exists a point such that then from (5.4) it follows that since is even. Therefore we may assume that for all
Following the proof of Theorem 1.2, we note Lemmas 3.2 and 3.3 continue to hold without any changes with the same definition for as before.
We assume and obtain a contradiction. Since using (5.4) we conclude
This, combined with strict convexity of gives a contradiction as in case (i) of Theorem 1.2. ∎
Proof of Proposition 1.8. We may assume that and denote the (not necessarily distinct) principal curvatures of For the proof, it is sufficient to replace the function appearing in the original proof ([11], Theorem ) by the function
Suppose for all Setting and using we compute
In the third line of the previous computation, we have used the formulas ([6], proposition 3.2)
and
to compute
Assume The definition of gives
Let and Then
Since is constant using the above we obtain
Noting we conclude
It follows that
for some constant We note that this relation is also valid for Rest of the argument is the same as in [11].